For some NONEMPTY independent events A and B we have ‏Pr(A) = Pr(B) = Pr(AnB),then


Select one:

A. Event A is impossible and event B is certain

B. Both A and B are impossible events

C. Both A and B are certain events

D. Event A is certain and event B is impossible


1
Expert's answer
2021-06-14T16:05:46-0400

Solution:

Since, A and B are independent events, then

"P(A \cap B)=P(A).P(B)
\\=P(A).P(A)=[P(A)]^2"

But it is given that "P(A \cap B)=P(A)"

Then, "[P(A)]^2=P(A)"

"\Rightarrow [P(A)]^2-P(A)=0
\\\Rightarrow P(A)[P(A)-1]=0
\\ \Rightarrow P(A)=0 \ or \ P(A)=1"

If "P(A)=0 \ then \ P(B)=0"

And if "P(A)=1 \ then \ P(B)=1"

Thus, all 4 options are true.